# Determining unit groups and $\mathrm{K}_{1}$ of finite rings

### v1

@@ 1. Introduction
Limitations of such an undertaking are immediate once one considers the situation for commutative rings, for which computations with and in unit groups are a classical topic in algorithmic number theory and cryptography. For example, determining the order $\lvert(\mathbf{Z}/n\mathbf{Z})^{\times}\rvert$ of the unit group of the finite ring $\mathbf{Z}/n\mathbf{Z}$ is equivalent to the problem of factoring $n$; see [Mil76, Sho09]. For a finite field $\mathbf{F}_{q}$, determining a generator of the unit group amounts to finding a primitive root (which reduces to factoring $q-1$), while expressing units with respect to a generator is the discrete logarithm problem for finite fields. Given these observations, it is reasonable to expect that determining unit groups of arbitrary finite rings is related to these “number theoretic obstacles”, the problems of factoring integers (Factor) and of computing discrete logarithms in finite fields (DiscLog).
@@ 2.1. Abelian groups and finite rings
    Any finite abelian group $A=\mathbf{Z}/n_{1}\mathbf{Z}\times\dotsb\times\mathbf{Z}/n_{r}\mathbf{Z}$ can be regarded as a black box group. The group $A$ is isomorphic to a finitely presented group with $r$ generators and $r+r(r-1)/2$ relators,
@@
    Let $G=\mathbf{F}_{q}^{\times}$ be the multiplicative group of a finite field of order $q$, which is cyclic of order $q-1$ and thus isomorphic to $\langle x\,|\,x^{q-1}\rangle$. Turning this into an effective presentation reduces (by definition) in polynomial time to PrimitiveRoot and DiscLog.
@@
We may assume that $n\geq 2$. The field $k$, a finite extension of some prime field $\mathbf{F}_{p}$ of degree $m$, is encoded as a simple extension $\mathbf{F}_{p}(\alpha)$ via a minimal polynomial in $\mathbf{F}_{p}[T]$, and thus of length $m\cdot\log(p)=\log(\lvert k\rvert)$. For $i,j\in\{1,\dotsc,n\}$ with $i\neq j$ and $\beta\in k$ denote by $e_{ij}(\beta)$ the elementary matrix with $1$’s on the diagonal and $\beta$ at position $(i,j)$. Since $k$ is a field, generated as an abelian group by $1,\alpha,\dotsc,\alpha^{m-1}$, $\operatorname{SL}_{n}(k)$ is generated by the elementary matrices $e_{ij}(\alpha^{r})$, $i\neq j$, $0\leq r<m$. Also, by applying row and column operations, there exists a polynomial-time algorithm that expresses an element $\operatorname{SL}_{n}(k)$ as a word in these generators.
@@ 4. Presentations and generators of unit groups
$1\to 1+I_{i-1}\to 1+I_{i}\to(1+I_{i})/(1+I_{i-1})\to 1$
Note that $1+I_{0}=\{1\}$ since $I^{2^{k}}=\{0\}$. By Lemma 4.1 for each $0\leq i\leq k$ we can determine effective presentations of $(1+I_{i})/(1+I_{i-1})$ in polynomial time and these presentations have at most $r$ generators and $r^{2}$ relators. Applying Corollary 3.3 shows that we can determine an effective presentation of $1+I$ in polynomial time. Note that as $n\leq\log_{2}(\lvert R\rvert)$ (the maximal length of a series of proper subgroups), $n$ is polynomial in the length of $R$ and the claim about the complexity follows. ∎
@@ 4.2. The semisimple case
    the problem of determining generators of the unit group reduces in probabilistic polynomial time to PrimitiveRoot.
@@
    the problem of determining generators of unit groups reduces in probabilistic polynomial time to Factor.
@@ 5.1. Abelianization
For a group $G$, denote by $G^{\mathrm{ab}}=G/[G,G]$ the abelianization (or maximal abelian quotient) of $G$. For a finite ring $R$, we consider the problem of computing $R^{\times\mathrm{ab}}$, a problem which we first have to make precise. Considering $R^{\times}$ as a black box group, by determining $R^{\times\mathrm{ab}}$ we mean determining a finite abelian group $A$ together with a surjective homomorphism $R^{\times}\to A$ with kernel $[R^{\times},R^{\times}]$ and for which images and preimages can be determined in probabilistic polynomial time (in particular $A\cong R^{\times\mathrm{ab}}$).
@@ 6. Relations between computational problems
    computing $\operatorname{K_{1}}1$ of finite rings,
@@ 8.1. First $K$-group for modular group rings of small groups
In [Mag06], Magurn determines $\operatorname{K_{1}}1(\mathbf{F}_{p}[G])$ for different finite groups $G$, including dihedral groups of order $2m$, $m=2^{r}\geq 4$ (a result which was stated in Keating [Kea82]), the alternating groups $\mathfrak{A}_{4}$, $\mathfrak{A}_{5}$, the symmetric group $\mathfrak{S}_{4}$, and for arbitrary groups of the form $\mathrm{C}_{2}^{m}\times G$. We have used our algorithm to determine $\mathbf{F}_{2}[G]^{\times\mathrm{ab}}$ and $\operatorname{K_{1}}1(\mathbf{F}_{2}[G])$ for all nonabelian groups of even order up to $32$. (We have also computed $\mathbf{F}_{p}[G]^{\times}$, but there is no convenient way to present the results.) The results are presented in Table 8.1. For a finite group $G$ with ID $(i,j)$ in the library of small groups [BEO02], we display unique invariants of the finite abelian groups $\mathbf{F}_{2}[G]^{\times\mathrm{ab}}$ and $\mathbf{F}_{2}[G]^{\times\mathrm{ab}}/\operatorname{K_{1}}1(\mathbf{F}_{2}[G])$. A group $(\mathbf{Z}/d_{1}\mathbf{Z})^{n_{1}}\times\dotsb\times(\mathbf{Z}/d_{r}\mathbf{Z})^{n_{r}}$ with $d_{1}\mid d_{2}\dotsm\mid d_{r}$ is represented by a list $[d_{1}^{n_{1}},\dotsc,d_{r}^{n_{r}}]$ of positive integers together with exponents, where we omit $n_{i}$ if the value is equal to $1$. The results are consistent with
@@
$\mathfrak{S}_{3}$ $(6,1)$ $[2^{2}]$ $[2]$
$\mathrm{D}_{8}$ $(8,3)$ $[2^{2},4]$ [][][ ]
$\mathrm{Q}_{8}$ $(8,4)$ $[2^{2},4]$ [][][ ]
$\mathrm{D}_{10}$ $(10,1)$ $[6]$ [][][ ]
$\mathrm{Q}_{12}$ $(12,1)$ $[2^{3},4]$ $[2]$
$\mathfrak{A}_{4}$ $(12,3)$ $[6]$ [][][ ]
$\mathrm{D}_{12}$ $(12,4)$ $[2^{5}]$ $[2]$
$\mathrm{D}_{14}$ $(14,1)$ $[14]$ [][][ ]
$\mathrm{C}_{2}^{2}\rtimes\mathrm{C}_{4}$ $(16,3)$ $[2^{5},4^{2}]$ [][][ ]
$\mathrm{C}_{4}\rtimes\mathrm{C}_{4}$ $(16,4)$ $[2^{5},4^{2}]$ [][][ ]
$\mathrm{M}_{4}(2)$ $(16,6)$ $[2^{3},4^{3}]$ [][][ ]
$\mathrm{D}_{16}$ $(16,7)$ $[2^{3},8]$ [][][ ]
$\mathrm{SD}_{16}$ $(16,8)$ $[2^{3},8]$ [][][ ]
$\mathrm{Q}_{16}$ $(16,9)$ $[2^{3},8]$ [][][ ]
$\mathrm{C}_{2}\times\mathrm{D}_{8}$ $(16,11)$ $[2^{7},4]$ [][][ ]
$\mathrm{C}_{2}\times\mathrm{Q}_{8}$ $(16,12)$ $[2^{7},4]$ [][][ ]
$\mathrm{C}_{4}\circ\mathrm{D}_{8}$ $(16,13)$ $[2^{7},4]$ [][][ ]
@@
$\mathrm{\mathrm{C}_{3}}\times\mathfrak{S}_{3}$ $(18,3)$ $[2^{2},6^{2}]$ $[2]$
$\mathrm{C}_{3}\rtimes\mathfrak{S}_{3}$ $(18,4)$ $[2^{5}]$ $[2^{4}]$
$\mathrm{Q}_{20}$ $(20,1)$ $[2^{3},12]$ [][][ ]
$\mathrm{F}_{5}$ $(20,3)$ $[2,4]$ [][][ ]
$\mathrm{D}_{20}$ $(20,4)$ $[2^{4},6]$ [][][ ]
$\mathrm{D}_{22}$ $(22,1)$ $[62]$ [][][ ]
$\mathrm{C}_{3}\rtimes\mathrm{C}_{8}$ $(24,1)$ $[2^{4},4^{2},8]$ $[2]$
$\operatorname{SL}_{2}(\mathbf{F}_{3})$ $(24,3)$ $[2^{2},12]$ [][][ ]
$\mathrm{Q}_{24}$ $(24,4)$ $[2^{4},4^{2}]$ $[2]$
$\mathrm{C}_{4}\times\mathfrak{S}_{3}$ $(24,5)$ $[2^{7},4^{2}]$ $[2]$
$\mathrm{D}_{24}$ $(24,6)$ $[2^{4},4^{2}]$ $[2]$
$\mathrm{C}_{2}\times\mathrm{Q}_{12}$ $(24,7)$ $[2^{9},4]$ $[2]$
$(24,8)$ $[2^{6},4]$ $[2]$
$\mathrm{C}_{3}\times\mathrm{D}_{8}$ $(24,10)$ $[2^{6},4^{2},12]$ [][][ ]
$\mathrm{C}_{3}\times\mathrm{Q}_{8}$ $(24,11)$ $[2^{6},4^{2},12]$ [][][ ]
$\mathfrak{S}_{4}$ $(24,12)$ $[2^{2},4]$ $[2]$
$\mathrm{C}_{2}\times\mathfrak{A}_{4}$ $(24,13)$ $[2^{4},6]$ [][][ ]
$\mathrm{C}_{2}^{2}\times\mathfrak{S}_{3}$ $(24,14)$ $[2^{11}]$ $[2]$
@@
$\mathrm{Q}_{28}$ $(28,1)$ $[2^{4},28]$ [][][ ]
$\mathrm{D}_{28}$ $(28,3)$ $[2^{5},14]$ [][][ ]
$\mathrm{C}_{5}\times\mathfrak{S}_{3}$ $(30,1)$ $[2^{4},30^{2}]$ $[2]$
$\mathrm{C}_{3}\times\mathrm{D}_{10}$ $(30,2)$ $[3,6^{3}]$ [][][ ]
@@
$(32,2)$ $[2^{13},4^{3}]$ [][][ ]
$(32,4)$ $[2^{7},4^{6}]$ [][][ ]
$\mathrm{C}_{2}^{2}\rtimes C_{8}$ $(32,5)$ $[2^{10},4^{3},8]$ [][][ ]
$(32,6)$ $[2^{4},4^{3}]$ [][][ ]
$(32,7)$ $[2^{4},4^{3}]$ [][][ ]
$(32,8)$ $[2^{4},4^{3}]$ [][][ ]
$(32,9)$ $[2^{8},4,8]$ [][][ ]
$(32,10)$ $[2^{8},4,8]$ [][][ ]
$\mathrm{C}_{4}\wr C_{2}$ $(32,11)$ $[2^{6},4^{2},8]$ [][][ ]
$\mathrm{C}_{4}\rtimes C_{8}$ $(32,12)$ $[2^{10},4^{3},8]$ [][][ ]
$(32,13)$ $[2^{8},4,8]$ [][][ ]
$(32,14)$ $[2^{8},4,8]$ [][][ ]
$(32,15)$ $[2^{6},4^{2},8]$ [][][ ]
$M_{5}(2)$ $(32,17)$ $[2^{7},4^{3},8^{2}]$ [][][ ]
$\mathrm{D}_{32}$ $(32,18)$ $[2^{4},4,16]$ [][][ ]
$\mathrm{SD}_{32}$ $(32,19)$ $[2^{4},4,16]$ [][][ ]
$\mathrm{Q}_{32}$ $(32,20)$ $[2^{4},4,16]$ [][][ ]
$(32,22)$ $[2^{15},4^{2}]$ [][][ ]
$(32,23)$ $[2^{15},4^{2}]$ [][][ ]
$(32,24)$ $[2^{13},4^{3}]$ [][][ ]
$\mathrm{C}_{4}\times D_{4}$ $(32,25)$ $[2^{13},4^{3}]$ [][][ ]
$\mathrm{C}_{4}\times Q_{8}$ $(32,26)$ $[2^{13},4^{3}]$ [][][ ]
$\mathrm{C}_{2}^{2}\wr C_{2}$ $(32,27)$ $[2^{7},4^{3}]$ [][][ ]
$(32,28)$ $[2^{9},4^{2}]$ [][][ ]
$(32,29)$ $[2^{9},4^{2}]$ [][][ ]
$(32,30)$ $[2^{7},4^{3}]$ [][][ ]
$(32,31)$ $[2^{7},4^{3}]$ [][][ ]
$(32,32)$ $[2^{7},4^{3}]$ [][][ ]
$(32,33)$ $[2^{7},4^{3}]$ [][][ ]
$(32,34)$ $[2^{7},4^{3}]$ [][][ ]
$(32,35)$ $[2^{7},4^{3}]$ [][][ ]
$\mathrm{C}_{2}\times M_{4}(2)$ $(32,37)$ $[2^{13},4^{3}]$ [][][ ]
$(32,38)$ $[2^{13},4^{3}]$ [][][ ]
$\mathrm{C}_{2}\times D_{8}$ $(32,39)$ $[2^{10},8]$ [][][ ]
$\mathrm{C}_{2}\times{\rm SD}_{16}$ $(32,40)$ $[2^{10},8]$ [][][ ]
$\mathrm{C}_{2}\times Q_{16}$ $(32,41)$ $[2^{10},8]$ [][][ ]
$(32,42)$ $[2^{10},8]$ [][][ ]
$(32,43)$ $[2^{7},8]$ [][][ ]
$(32,44)$ $[2^{7},8]$ [][][ ]
$\mathrm{C}_{2}^{2}\times D_{4}$ $(32,46)$ $[2^{17},4]$ [][][ ]
$\mathrm{C}_{2}^{2}\times Q_{8}$ $(32,47)$ $[2^{17},4]$ [][][ ]
$(32,48)$ $[2^{17},4]$ [][][ ]
$(32,49)$ $[2^{14},4]$ [][][ ]
$(32,50)$ $[2^{14},4]$ [][][ ]

### v2

@@ 1. Introduction
Limitations of such an undertaking are immediate once one considers the situation for commutative rings, for which computations with and in unit groups are a classical topic in algorithmic number theory and cryptography. For example, determining the order $\lvert(\mathbf{Z}/n\mathbf{Z})^{\times}\rvert$ of the unit group of the finite ring $\mathbf{Z}/n\mathbf{Z}$ is equivalent to the problem of factoring $n$; see [Mil76, Sho09]. For a finite field $\mathbf{F}_{q}$, determining a generator of the unit group amounts to finding a primitive root (which reduces to factoring q1q-1), while expressing units with respect to a generator is the discrete logarithm problem for finite fields. Given these observations, it is reasonable to expect that determining unit groups of arbitrary finite rings is related to these “number theoretic obstacles”, the problems of factoring integers (Factor) and of computing discrete logarithms in finite fields (DiscLog).
@@ 2.1. Abelian groups and finite rings
    Any finite abelian group $A=\mathbf{Z}/n_{1}\mathbf{Z}\times\dotsb\times\mathbf{Z}/n_{r}\mathbf{Z}$ can be regarded as a black box group. The group $A$ is isomorphic to a finitely presented group with $r$ generators and r+r(r1)/2r+r(r-1)/2 relators,
@@
    Let $G=\mathbf{F}_{q}^{\times}$ be the multiplicative group of a finite field of order $q$, which is cyclic of order q1q-1 and thus isomorphic to $\langle x\,|\,x^{q-1}\rangle$. Turning this into an effective presentation reduces (by definition) in polynomial time to PrimitiveRoot and DiscLog.
@@
[…]
@@ 4. Presentations and generators of unit groups
$1\to 1+I_{i-1}\to 1+I_{i}\to(1+I_{i})/(1+I_{i-1})\to 1.$
[…]
@@ 4.2. The semisimple case
    the problem of determining generators of the unit group reduces in probabilistic polynomial time to PrimitiveRoot, and
@@
    the problem of determining generators of unit groups reduces in probabilistic polynomial time to Factor, and
@@ 5.1. Abelianization
For a group $G$, denote by $G^{\mathrm{ab}}=G/[G,G]$ the abelianization (or maximal abelian quotient) of $G$. For a finite ring $R$, we consider the problem of computing $R^{\times\mathrm{ab}}$, a problem which we first have to make precise. Considering $R^{\times}$ as a black box group, by determining $R^{\times\mathrm{ab}}$ we mean determining a finite abelian group $A$ together with a surjective homomorphism $R^{\times}\to A$ with kernel $R^{\times},R^{\times}]$ and for which images and preimages can be determined in probabilistic polynomial time (in particular $A\cong R^{\times\mathrm{ab}}$).
@@ 6. Relations between computational problems
    computing $\operatorname{K_{1}}1$ of finite rings.
@@ 8.1. First K-group for modular group rings of small groups
In [Mag06], Magurn determines $\operatorname{K_{1}}1(\mathbf{F}_{p}[G])$ for different finite groups $G$, including dihedral groups of order $2m$, $m=2^{r}\geq 4$ (a result which was stated in Keating [Kea82]), the alternating groups $\mathfrak{A}_{4}$, $\mathfrak{A}_{5}$, the symmetric group $\mathfrak{S}_{4}$, and for arbitrary groups of the form $\mathrm{C}_{2}^{m}\times G$. We have used our algorithm to determine $\mathbf{F}_{2}[G]^{\times\mathrm{ab}}$ and $\operatorname{K_{1}}1(\mathbf{F}_{2}[G])$ for all nonabelian groups of even order up to $32$. (We have also computed $\mathbf{F}_{p}[G]^{\times}$, but there is no convenient way to present the results.) The results are presented in Table 1. For a finite group $G$ with ID $(i,j)$ in the library of small groups [BEO02], we display unique invariants of the finite abelian groups $\mathbf{F}_{2}[G]^{\times\mathrm{ab}}$ and $\mathbf{F}_{2}[G]^{\times\mathrm{ab}}/\operatorname{K_{1}}1(\mathbf{F}_{2}[G])$. A group $(\mathbf{Z}/d_{1}\mathbf{Z})^{n_{1}}\times\dotsb\times(\mathbf{Z}/d_{r}\mathbf{Z})^{n_{r}}$ with $d_{1}\mid d_{2}\dotsm\mid d_{r}$ is represented by a list $d_{1}^{n_{1}},\dotsc,d_{r}^{n_{r}}]$ of positive integers together with exponents, where we omit $n_{i}$ if the value is equal to $1$. The results are consistent with
@@
Table 1. First K-group for modular group rings of small groups.
@@
$\mathfrak{S}_{3}$ $(6,1)$ $^{2}] [2][2$]
$\mathrm{D}_{8}$ $(8,3)$ $2^{2},4] [$][]
$\mathrm{Q}_{8}$ $(8,4)$ $2^{2},4] [$][]
$\mathrm{D}_{10}$ $(10,1)$ $[6]$ $[]$
$\mathrm{Q}_{12}$ $(12,1)$ $^{3},4] [2][2$]
$\mathfrak{A}_{4}$ $(12,3)$ $[6]$ $[]$
$\mathrm{D}_{12}$ $(12,4)$ $^{5}] [2][2$]
$\mathrm{D}_{14}$ $(14,1)$ $[14]$ $[]$
$\mathrm{C}_{2}^{2}\rtimes\mathrm{C}_{4}$ $(16,3)$ $2^{5},4^{2}] [$][]
$\mathrm{C}_{4}\rtimes\mathrm{C}_{4}$ $(16,4)$ $2^{5},4^{2}] [$][]
$\mathrm{M}_{4}(2)$ $(16,6)$ $2^{3},4^{3}] [$][]
$\mathrm{D}_{16}$ $(16,7)$ $2^{3},8] [$][]
$\mathrm{SD}_{16}$ $(16,8)$ $2^{3},8] [$][]
$\mathrm{Q}_{16}$ $(16,9)$ $2^{3},8] [$][]
$\mathrm{C}_{2}\times\mathrm{D}_{8}$ $(16,11)$ $2^{7},4] [$][]
$\mathrm{C}_{2}\times\mathrm{Q}_{8}$ $(16,12)$ $2^{7},4] [$][]
$\mathrm{C}_{4}\circ\mathrm{D}_{8}$ $(16,13)$ $2^{7},4] [$][]
@@
$\mathrm{\mathrm{C}_{3}}\times\mathfrak{S}_{3}$ $(18,3)$ $^{2},6^{2}] [2][2$]
$\mathrm{C}_{3}\rtimes\mathfrak{S}_{3}$ $(18,4)$ $2^{5}] [$$2^{4}]$
$\mathrm{Q}_{20}$ $(20,1)$ $2^{3},12] [$][]
$\mathrm{F}_{5}$ $(20,3)$ $[2,4]$ $[]$
$\mathrm{D}_{20}$ $(20,4)$ $2^{4},6] [$][]
$\mathrm{D}_{22}$ $(22,1)$ $[62]$ $[]$
$\mathrm{C}_{3}\rtimes\mathrm{C}_{8}$ $(24,1)$ $^{4},4^{2},8] [2][2$]
$\operatorname{SL}_{2}(\mathbf{F}_{3})$ $(24,3)$ $2^{2},12] [$][]
$\mathrm{Q}_{24}$ $(24,4)$ $^{4},4^{2}] [2][2$]
$\mathrm{C}_{4}\times\mathfrak{S}_{3}$ $(24,5)$ $^{7},4^{2}] [2][2$]
$\mathrm{D}_{24}$ $(24,6)$ $^{4},4^{2}] [2][2$]
$\mathrm{C}_{2}\times\mathrm{Q}_{12}$ $(24,7)$ $^{9},4] [2][2$]
$(24,8)$ $^{6},4] [2][2$]
$\mathrm{C}_{3}\times\mathrm{D}_{8}$ $(24,10)$ $2^{6},4^{2},12] [$][]
$\mathrm{C}_{3}\times\mathrm{Q}_{8}$ $(24,11)$ $2^{6},4^{2},12] [$][]
$\mathfrak{S}_{4}$ $(24,12)$ $^{2},4] [2][2$]
$\mathrm{C}_{2}\times\mathfrak{A}_{4}$ $(24,13)$ $2^{4},6] [$][]
$\mathrm{C}_{2}^{2}\times\mathfrak{S}_{3}$ $(24,14)$ $^{11}] [2][2$]
@@
$\mathrm{Q}_{28}$ $(28,1)$ $2^{4},28] [$][]
$\mathrm{D}_{28}$ $(28,3)$ $2^{5},14] [$][]
$\mathrm{C}_{5}\times\mathfrak{S}_{3}$ $(30,1)$ $^{4},30^{2}] [2][2$]
$\mathrm{C}_{3}\times\mathrm{D}_{10}$ $(30,2)$ $3,6^{3}] [$][]
@@
$(32,2)$ $2^{13},4^{3}] [$][]
$(32,4)$ $2^{7},4^{6}] [$][]
$\mathrm{C}_{2}^{2}\rtimes\mathrm{C}_{8}$ $(32,5)$ $2^{10},4^{3},8] [$][]
$(32,6)$ $2^{4},4^{3}] [$][]
$(32,7)$ $2^{4},4^{3}] [$][]
$(32,8)$ $2^{4},4^{3}] [$][]
$(32,9)$ $2^{8},4,8] [$][]
$(32,10)$ $2^{8},4,8] [$][]
$\mathrm{C}_{4}\wr\mathrm{C}_{2}$ $(32,11)$ $2^{6},4^{2},8] [$][]
$\mathrm{C}_{4}\rtimes\mathrm{C}_{8}$ $(32,12)$ $2^{10},4^{3},8] [$][]
$(32,13)$ $2^{8},4,8] [$][]
$(32,14)$ $2^{8},4,8] [$][]
$(32,15)$ $2^{6},4^{2},8] [$][]
$\mathrm{M}_{5}(2)$ $(32,17)$ $2^{7},4^{3},8^{2}] [$][]
$\mathrm{D}_{32}$ $(32,18)$ $2^{4},4,16] [$][]
$\mathrm{SD}_{32}$ $(32,19)$ $2^{4},4,16] [$][]
$\mathrm{Q}_{32}$ $(32,20)$ $2^{4},4,16] [$][]
$(32,22)$ $2^{15},4^{2}] [$][]
$(32,23)$ $2^{15},4^{2}] [$][]
$(32,24)$ $2^{13},4^{3}] [$][]
$\mathrm{C}_{4}\times\mathrm{D}_{4}$ $(32,25)$ $2^{13},4^{3}] [$][]
$\mathrm{C}_{4}\times\mathrm{Q}_{8}$ $(32,26)$ $2^{13},4^{3}] [$][]
$\mathrm{C}_{2}^{2}\wr\mathrm{C}_{2}$ $(32,27)$ $2^{7},4^{3}] [$][]
$(32,28)$ $2^{9},4^{2}] [$][]
$(32,29)$ $2^{9},4^{2}] [$][]
$(32,30)$ $2^{7},4^{3}] [$][]
$(32,31)$ $2^{7},4^{3}] [$][]
$(32,32)$ $2^{7},4^{3}] [$][]
$(32,33)$ $2^{7},4^{3}] [$][]
$(32,34)$ $2^{7},4^{3}] [$][]
$(32,35)$ $2^{7},4^{3}] [$][]
$\mathrm{C}_{2}\times\mathrm{M}_{4}(2)$ $(32,37)$ $2^{13},4^{3}] [$][]
$(32,38)$ $2^{13},4^{3}] [$][]
$\mathrm{C}_{2}\times\mathrm{D}_{8}$ $(32,39)$ $2^{10},8] [$][]
$\mathrm{C}_{2}\times{\rm SD}_{16}$ $(32,40)$ $2^{10},8] [$][]
$\mathrm{C}_{2}\times\mathrm{Q}_{16}$ $(32,41)$ $2^{10},8] [$][]
$(32,42)$ $2^{10},8] [$][]
$(32,43)$ $2^{7},8] [$][]
$(32,44)$ $2^{7},8] [$][]
$\mathrm{C}_{2}^{2}\times\mathrm{D}_{4}$ $(32,46)$ $2^{17},4] [$][]
$\mathrm{C}_{2}^{2}\times\mathrm{Q}_{8}$ $(32,47)$ $2^{17},4] [$][]
$(32,48)$ $2^{17},4] [$][]
$(32,49)$ $2^{14},4] [$][]
$(32,50)$ $2^{14},4] [$][]